# A systematic comparison of measures for k-anonymity in networks

### v1

@@ 6.3 Comparing runtimes
As discussed in Section 2.3, a second important aspect of choosing a measure is whether it is feasible to compute for a network of a given size. It is worth nothing that computing anonymity consists of two parts: computing the value for a node using the chosen measure and book-keeping operations to determine, based on the values computed, the right equivalence class for each node. In Figure 4, we show the runtime for each of the measures on the set of considered real-world networks. We first focus on the results for $d=1$ in the top figure.
@@
Overall, degree and vrq are least expensive to compute. An explanation for this is that for all other measures the $d$-neighborhood of the target nodes need to be extracted, which is relatively time consuming. Instead, for degree merely the degree of the considered node needs to be determined, and for vrq one only needs to iterate over the nodes at a certain distance and determine their degree. For the neighborhood based measures, computing the degree distribution as required for degdist or determining the canonical labeling as required for $d$-$k$-anonymity and hybrid, can consume a noticeable amount of the total runtime. For most networks count is less time consuming than the other neighborhood based measures, but for a few other networks, such as “Radoslaw emails" and “FB Wellesley22" more time appears to be consumed by book-keeping operations. As a result, count achieves runtimes similar to or higher than other measures, including $d$-$k$-anonymity and hybrid.
@@
When moving to the results for $d=2$, shown in the bottom of Figure 4, we see that the runtimes are more diverse, up to  15 minutes for some of the denser networks. For several large networks, “The marker cafe" and “Anybeat", the values could not be computed for all measures within the set time limit. Here we see again that for some networks the runtime is much larger for count. This is likely due to the lower uniqueness achieved by this measure for $d=1$. As a result, the value for more nodes should be determined and more book-keeping operations are required to obtain the equivalence partition for $d=2$. It appears that hybrid is often the most time consuming measure as it combines vrq with $d$-$k$-anonymity.
6.4 Anonymity-cascade
The results from the previous section showed that measures that reach further achieve a much higher uniqueness. In this section, we investigate the effect on uniqueness when using a different method to increase the reach even further, namely by cascading the information of unique nodes to identify more nodes. To do so, we measure uniqueness using a generic variant of anonymity-cascade [12], which is explained in Section 4.3.5. The algorithm consists of two steps. First, we identify nodes using the initial measure. Second, we use the cascade measure to identify more nodes in a cascading fashion. This is also illustrated in Figure 5. The cascading process can continue for multiple levels. If only one level of cascading is used, we refer to this as $C_{1}$. The process can then repeated until no more nodes can be identified, referred to as cascading final $C_{f}$.
The combinations of initial and cascade measure consist of all 36 combinations of the six measures in Table 3. This allows us to investigate all combinations of all attacker knowledge. Besides the cases where the levels of knowledge are equal, this also includes the cases where an attacker has little starting knowledge and more local knowledge surrounding the identified nodes, hence more cascade knowledge. We also include scenarios where an attacker has a high level of knowledge on some nodes, which could for example be the high degree nodes, but little knowledge on the surrounding nodes. While we fully acknowledge that not every combination of measures is equally realistic, we chose to include all to observe the cascading effect of further reach on the uniqueness.
Figure 6: Uniqueness using anonymity-cascade one level. Each figure corresponds to a different initial measure (triangle and line) and is combined with all of the cascade measures (dots). The grey bar indicates the highest obtained uniqueness for the network given the initial measure. Measures used are: degree (lightblue), count (blue), degdist (dark blue), $d$-$k$-anonymity (black triangle), vrq (red) and hybrid (pink).
We focus on one level of cascading, as this is a more realistic scenario. The results for cascading final can be found in Appendix C. Figure 6 shows the uniqueness for anonymity-cascade with one level of cascading. Each subfigure shows results for a different initial measure, denoted by the line with triangle, and results of one level of cascading, denoted by the colored dots using each of the six measures as cascading measure. The grey bar is added for readability of the results and shows the highest uniqueness obtained.
First, turning to the results for degree, we see a very low initial uniqueness. However, one level of cascading overall has a large impact on uniqueness. When using degree as cascading measure, this impact is smaller in several cases. In most cases using vrq as cascading measure has the largest impact on the uniqueness.
The results of degree however, still differ a lot from results for the other measures. Results show that count, degdist and $d$-$k$-anonymity achieve similar results, where using vrq as cascade measure does increase uniqueness. With vrq as initial measure we do observe both a higher initial uniqueness and uniqueness after cascading. Turning to the results for cascading final in Appendix C, we see a large difference for degree and count. For degdist and $d$-$k$-anonymity there is some visible increase, for vrq there is no large increase compared to one level of cascading.
Overall, these results show, similar to the results of Section 6.2, that by reaching further, with one or more cascading steps, the uniqueness increases immensely. This effect is especially clear when using degree as initial measure.
@@ 7 Conclusions
In this paper, we have investigated the problem of privacy aware sharing of network data. First, we introduced the aspects of the problem that a user should take into account when choosing an approach for achieving anonymity in networks, being data utility, privacy and runtime. Then, we provided an overview of the methods introduced for achieving privacy in networks and related these to the different aspects. The approaches can be grouped based on the type of output they produce: an interactive model, synthetic data, an intermediate representation or a perturbed version of the network. Finally, as a particularly suitable approach for privacy aware sharing of network data, we focus on $k$-anonymity, in which the choice of measure models the attacker scenario to protect against and hence is important for a user to take into account. Therefore, we provide an overview of the different measures introduced in literature resulting in a categorization into degree based, neighborhood based and automorphism based measures. Next, we compared the selected $k$-anonymity measures theoretically, showing how they can be distinguished based on how far they reach, and the completeness of the structural information they take into account. This resulted in an ordering of the measures based on strictness.
@@
To better understand how the different measures impact the anonymity, we compared the uniqueness obtained by different measures on a wide range of empirical networks. The measure of degree often finds only few unique nodes and shows to be a very ineffective approach for identifying nodes. If an attacker has additional information, such as the degree of the directly neighboring nodes (vrq), or the number of triangles in the 1-neighborhood (count), they are likely able to uniquely identify many more nodes, often close to the number that can be identified by $d$-$k$-anonymity. Most importantly, the results showed that it is more de-anonymizing to have incomplete information that accounts for structural information that reaches further, beyond the direct neighborhood of the considered node, than complete information concerning its direct neighborhood. Even more so, by means of experiments using the hybrid measure we show that often having (more) complete information only has a small additional effect on the uniqueness if the reach remains the same. At the same time these less complete measures, even if they reach further, require less time to compute anonymity. Hence, when choosing a measure for anonymity, it would be better to choose a measure that reaches beyond the direct neighborhood (assuming that this is a realistic attacker scenario), rather than a measure that accounts for as complete as possible structural information. Lastly, we investigate the effect of extending the reach of the anonymity measure, namely by reusing unique nodes in a cascading fashion to identify more nodes. Just one level of cascading has an immense impact on the anonymity. This result is especially large when using degree as starting measure, which identifies very few nodes.
There are still a lot of possibilities for future work. Firstly, most measures for anonymity so far focus on the direct neighborhood. At the same time our results on measuring anonymity show the necessity to look beyond the 1-neighborhood. A crucial next step would be to design anonymization algorithms accounting for measures that look beyond the 1-neighborhood of the node. Second, while this manuscript described how different measures affect certain aspects, some aspects are heavily influenced by the chosen anonymization algorithm. Therefore, another direction for future work, complementing this work on anonymity measurement would be to make a systematic overview of existing anonymization methods and how they affect runtime and utility.
Acknowledgments
This research was made possible by the Platform Digital Infrastructure SSH (http://www.pdi-ssh.nl). We would also like to thank the POPNET team (https://www.popnet.io) and the Leiden CNS group (https://www.computationalnetworkscience.org) for various helpful suggestions and discussions.
Appendix A Empirical differences between measures
@@
Appendix B Empirical differences and network properties
@@
Appendix C Anonymity-cascade

### v2

@@ 6.5 Comparing runtimes
In this section, we investigate the computational aspect by measuring the time required to compute uniqueness using each of the six anonymity measures. It is worth nothing that computing anonymity consists of two parts: computing the value for a node using the chosen measure and bookkeeping operations to determine, based on the values computed, the right equivalence class for each node to arrive at the network’s uniqueness value. In Figure 6, we show the runtime for each of the measures on the set of considered real-world networks. We first focus on the results for $d=1$ in the top figure.
Figure 6: Runtime in seconds (vertical axis) on different network datasets (horizontal axis) for different $k$-anonymity measures: degree (lightblue circle), count (blue square), degdist (dark blue circle), $d$-$k$-anonymity (black triangle), vrq (red circle) and hybrid (pink square).
@@
Overall, degree and vrq are the least expensive to compute. This can be explained by the fact that for the neighborhood based measures, count, degdist and $d$-$k$-anonymity, the $d$-neighborhood of the considered node needs to be extracted, which is relatively time consuming. Instead, for degree merely the degree of the considered node needs to be determined, and for vrq one only needs to iterate over the neighboring nodes and determine their degree. For the neighborhood based measures, computing the degree distribution as required for degdist or determining the canonical labeling as required for $d$-$k$-anonymity and hybrid, increases the runtime notably for some networks. For most networks count is less time consuming than the other neighborhood based measures except for a few other networks, such as “Radoslaw emails” and “FB Wellesley22”, for which more time appears to be consumed by bookkeeping operations. As a result, count achieves runtimes similar to or higher than other measures, including $d$-$k$-anonymity and hybrid.
@@
For $d=2$, shown in the bottom of Figure 6, we see that the runtimes are more diverse, up to  15 minutes for some of the denser networks. For several large networks, “The marker cafe” and “Anybeat”, the values could not be computed for $d$-$k$-anonymity and hybrid within the given time limit. This is due to the canonical labeling computation which is in very few cases expensive to compute [12] and can immensely increase the runtime. This unpredictability in runtime is not a desirable property. For $d=2$ we again notice that for some networks the runtime is much larger for count. This is likely due to the lower uniqueness achieved by this measure for $d=1$. As a result, the value for more nodes should be determined and more bookkeeping operations are required to obtain the equivalence partition for $d=2$. It appears that hybrid is often the most time consuming measure as it combines vrq and $d$-$k$-anonymity.
@@ 7 Conclusions
In this paper, we have systematically compared six measures for $k$-anonymity, both theoretically and empirically. The choice in anonymity measure strongly affects three important aspects, being, 1) the considered attacker scenario and its effectiveness, 2) the anonymity vs. utility trade-off, and 3) computation time.
First, we provided an overview of the different measures introduced in literature resulting in a categorization into degree based, neighborhood based and automorphism based measures. Then, we compared the selected $k$-anonymity measures theoretically, showing how they can be distinguished based on how far they reach, and the completeness of the structural information they take into account. Based on this, we derived a formal distance-parameterized strictness-based ordering of the measures. If an anonymity measure is more strict than another measure, this has three implications: 1) The more strict measure protects against all attacker scenarios the less strict measure protects against, 2) the anonymity of a network can be at most equal to that measured by the less strict measure, and 3) anonymization for the more strict measure requires at least as many graph alterations.
@@
To better understand how the different measures impact the three aspects, we performed four experiments on a wide range of empirical networks. First, we measured uniqueness of the networks to assess the effectiveness of the considered attacker scenario mimicked by the anonymity measure. Most importantly, the results showed that it is more de-anonymizing to have incomplete information that accounts for structural information that reaches further, beyond the direct neighborhood of the considered node, than complete information concerning its direct neighborhood. Even more so, by means of the hybrid measure we show that often having (more) complete information only has a small additional effect on the uniqueness if the reach remains the same. At the same time these less complete measures, even if they reach further, require less time to compute anonymity. Hence, when choosing a measure for anonymity, it would be better to choose a measure that reaches beyond the direct neighborhood (assuming that this is a realistic attacker scenario), rather than a measure that accounts for as complete as possible structural information.
Second, we investigate the effect of extending the reach of the anonymity measure by reusing unique nodes in a cascading fashion to identify more nodes. Just one level of cascading has an immense impact on the anonymity, especially when using degree as initial measure, which by itself identifies very few nodes.
Third, we investigated empirically how the chosen measure affects the trade-off between anonymity and utility by measuring the difference in performance on three downstream tasks during anonymization. The results showed that using more strict measures, which protect against more strict attacker scenarios, results in higher observed uniqueness and substantially more utility has to be sacrificed to obtain more anonymity. The count measure resulted in slightly better trade-offs than degdist and $d$-$k$-anonymity, which achieved similar results in terms of uniqueness and utility.
Lastly, we investigated the runtime required to measure anonymity with the different measures. The most affecting factor was the reach of the measures. As the neighborhood based measures, count, degdist and $d$-$k$-anonymity, need to extract the neighborhood of considered nodes, these measures overall required a longer runtime compared to degree and vrq which need solely to compute the degree of nodes.
There are still many possibilities for future work. Most measures for anonymity so far focus on the direct neighborhood. At the same time our results on measuring anonymity show the necessity to look beyond the 1-neighborhood. A crucial next step would be to design anonymization algorithms accounting for measures that look beyond the 1-neighborhood of the node. Furthermore, the existing anonymity measures can be extended to account for additional network properties, such as node or edge labels, or even temporality. Overall, this paper provides a foundation for future research on $k$-anonymity maximization algorithms of networks, for designing anonymity measures, and, in general, aiding in the selection of suitable existing measures in applied scenarios.
Appendix A Strictness
This appendix accompanies Section 5. Theorem 1 states a property of strictness: if a measure is more strict, anonymization through perturbation for this measure will require at least as many operations as for a less (or equally strict) measure. Theorem 2 and 3 are used to make the ordering in Figure 1.
For Theorem 1 we use the notion of graph alterations which is defined as operations that modify the graph, such as edge deletion and addition, and node deletion or addition.
Theorem 1.
Given a graph $G=(V,E)$ and measures $M_{1}\geq M_{2}$. To make $G$ anonymous for $M_{1}$, at least as many alterations of the graph are required as for $M_{2}$.
Proof.
From the notion of strictness (see Definition 1 in Section 5) it follows that if $G$ is anonymous for $M_{1}$, then $G$ is anonymous for $M_{2}$. This implies that if $G$ can be anonymized for $M_{1}$ with $X$ alterations, then it can also be anonymized for $M_{2}$ using the same $X$ alterations. Hence, to anonymize a graph for $M_{1}$, one needs at least as many edge deletions as for $M_{2}$. ∎
Theorem 2.
For $d\geq 1$, the following holds:
  • degree $\leq$ count$(d)$leq $(d)$leq $d$-$k$-anonymity $\leq$ hybrid$(d)$
  • vrq$(d)$leq hybrid$(d)$
For $d\geq 2$:
  • vrq$(d-1)$leq $d$-$k$-anonymity.
  • vrq$(d-1)$leq vrq$(d)$.
Proof.
Given a graph $G=(V,E)$ and nodes $v,w$:
  • degree $\leq$ count. Since the degree of the considered node equals the number of nodes in the $1$-neighborhood minus one, the degree can be derived from count at $d=1$. As a result, two nodes equivalent according to count are equivalent for degree, ensuring the requirement: $v\cong_{\textsc{count}}w\rightarrow v\cong_{\textsc{degree}}w$.
Theorem 3.
[…]
Proof.
We prove this theorem by use of the counterexamples in Figure 7. The figure shows one case for which count is able to distinguish between two nodes and vrq is not, an example showing the converse where vrq is able to distinguish between the nodes and count is not and an example where vrq is able to distinguish between two nodes and count$(d)$ and $d+1$ and is not. Hence, we can not conclude that one measure is more strict than the other. ∎
Appendix B Empirical differences between measures
@@
Appendix C Empirical differences and network properties
@@
Appendix D Anonymity-cascade
@@
Appendix E Measures and utility
This appendix accompanies Section 6.4 of the main paper containing the Pareto fronts found for 14 networks.
Figure 11: Pareto optimal solutions found in terms of uniqueness (horizontal axis) and performance on a downstream task (vertical axis, one per column) for 14 networks (rows). Each dot represents a solution found when anonymizing using edge sampling for one of the five anonymity measures (colors). Error bars indicate results $\pm$ one standard deviation.